OpenStudy (anonymous):

Solve the system of equations below by graphing them with a pencil and paper. Enter your answer as an ordered pair. y = 2x + 1 y = -x + 4

OpenStudy (anonymous):

Just plug in values and graph your answer

OpenStudy (anonymous):

Enter your answer as an ordered pair.

OpenStudy (anonymous):

What ladii's saying is put in numbers for your x, and find out what your y values are. Or, after plugging in 0 for x and getting y=2(0)+1 -> y= 0+ 1 -> y= 1, you can put your first dot on (0,1), and then using the slope of 2 from 2x, go 2 up, one right. Two up, one right. And then make the line following the pattern. Do the same procedure for y=-x+4. Plug in a zero, see where to put your first dot, and since the slope is NEGATIVE one, you go one down and to the right each time.

OpenStudy (anonymous):

I came up with this: http://www.scribblar.com/kuub1xb

OpenStudy (anonymous):

It's pretty rough but the best I can do on scribblar

OpenStudy (anonymous):

Yay scribblar! :D

OpenStudy (anonymous):

I just started using it today, still learning the ropes on scribblar:)

OpenStudy (anonymous):

i got 2,5